# EFlat-LoRA: Efficiently Seeking Flat Minima for Better Generalization in Fine-Tuning Large Language Models and Beyond

### v1

@@ Introduction
Flat-LoRA 95.1±0.5 54.4±1.3 86.4±0.8 82.7±1.0 82.7±1.2 96.7±0.2 83.1
EFlat-LoRA 91.9±1.7 54.7±1.6 85.7±0.7 82.1±0.6 82.8±0.2 96.8±0.4 82.3
Table 1: Experiments on few-shot RoBERTa (355M). Results marked with * are taken from (Li, Zhang, and He 2024).
To theoretically analyze the error of EFlat-LoRA, some necessary assumptions are listed below, all of which are common and standard when analyzing SAM optimization (Du et al. 2022) (Zhuang et al. 2022).
Assumption 1
@@
This convexity assumption is reasonable in the fine-tuning stage, as the model is typically close to a local minimum and the loss landscape is approximately convex in a local neighborhood (Malladi et al. 2023) (Kim, Kim, and Ryu 2025).
Theorem 2
[EMA pertubation approximate pertubation of SAM due to the convex of the loss landscape] Assume that during fine-tuning, the solution is already close to a local minimum and the local loss function is convex. Let the model weights at $i$-th iteration be $\mathbf{w}_{t}$. Under Assumptions 1, 2, and 3, let ${\rho_{t}}=\frac{{{\rho_{0}}}}{{\sqrt{t}}}$, the error between the sharpness calculated using the EMA perturbation ($S^{\text{EMA}}$) and that calculated using the original SAM perturbation ($S^{\text{SAM}}$) is bounded as follows:
$|\underbrace{\left[{L({w_{t}}+{{\hat{\varepsilon}}_{t-1}})-L({w_{t}})}\right]}_{S^{\text{EMA}}}-\underbrace{\left[{L({w_{t}}+{{\tilde{\varepsilon}}_{t}})-L({w_{t}})}\right]}_{S^{\text{SAM}}}|$ (19)
$\leq\left({\tau\frac{{{\rho_{0}}}}{{\sqrt{t-1}}}+G+{\sigma^{2}}}\right)\left({\frac{{{\rho_{0}}}}{{\sqrt{t}}}+{\rho_{0}}{{(1-\beta)}^{t-1}}+{\rho_{0}}}\right).$
@@
RoBERTa SST2 STS-B RTE QQP QNLI MRPC MNLI CoLA avg.$\uparrow$
@@
EFlat-LoRA 96.3±0.2 92.7±0.1 89.3±0.6 91.6±0.1 94.8±0.1 91.5±0.4 90.7±0.1 68.0±1.2 89.4
Table 2: Experiments on finetuning RoBERTa (355M). Results marked with $\dagger$ are taken from (Hu et al. 2022), and those with * are taken from (Li, Zhang, and He 2024).
@@
GPT-2 M (Flat-LoRA) 0.35M 69.2 8.72 46.6 71.5 2.51
GPT-2 M (EFlat-LoRA) 0.35M 69.7 8.77 46.6 71.7 2.53
@@
GPT-2 L (Flat-LoRA) 0.77M 70.0 8.83 46.8 71.8 2.53
GPT-2 L (EFlat-LoRA) 0.77M 70.2 8.84 46.8 71.8 2.52
Table 3: GPT-2 medium (M) and large (L) with different adaptation methods on the E2E NLG Challenge. Results marked with $\dagger$ are taken from (Hu et al. 2022).
@@
LoRA reduces the number of trainable parameters by decomposing weight updates as $\Delta\mathbf{W}\approx\mathbf{B}\mathbf{A}$, where $\mathbf{B}\in\mathbb{R}^{n\times r}$ and $\mathbf{A}\in\mathbb{R}^{r\times m}$ with $r\ll\min(n,m)$. Both Flat-LoRA and EFlat-LoRA retain this parameter efficiency:
$\text{P}_{\text{LoRA}}$ $=\text{P}_{\text{Flat-LoRA}}=\text{P}_{\text{EFlat-LoRA}}$ (20)
@@
However, Flat-LoRA and EFlat-LoRA introduce additional memory overhead. Specifically, Flat-LoRA temporarily stores the original values of $\mathbf{B}$ and $\mathbf{A}$, as well as the gradients of $\mathbf{A}$. The memory usage of Flat-LoRA is calibrated as follows:
@@
where $\text{M}_{\text{LoRA}}$ indicates the memory required by LoRA. The memory of EFlat-LoRA needs to maintain the EMA perturbation on $\mathbf{B}$ as follows:
$\text{M}_{\text{EFlat-LoRA}}=\text{M}_{\text{LoRA}}+O(2\times(nr+rm)).$ (22)
@@
For time complexity, suppose that the time complexity of optimizing the model with LoRA is $O(T)$, which mainly includes the time for forward and backward. Theoretically, the time complexity of Flat-LoRA is approximately as follows:
$\text{T}_{\text{Flat-LoRA}}\approx O(2T)=2\times\text{T}_{\text{LoRA}}.$ (23)
In contrast, the time complexity of EFlat-LoRA can be approximated as follows:
$\text{T}_{\text{EFlat-LoRA}}\approx O(T)=\text{T}_{\text{LoRA}}.$ (24)
@@
The best and second-best results are highlighted in bold and underline, respectively. Additional experimental details are provided in Section B of Supplementary file.
@@
Few-shot with RoBERTa-large. We first consider few-shot learning with EFlat-LoRA, aiming to fine-tune a language model using limited training data. Following the setup of (Li, Zhang, and He 2024), we adopt RoBERTa-large—a 355M-parameter language model—as the backbone. The results of the proposed EFlat-LoRA on RoBERTa-large are summarized in Table 1. The results show that Flat-LoRA outperforms all other methods with the highest average score (83.1), particularly excelling on SST-2, SNLI, and MNLI. EFlat-LoRA follows closely with an average score of 82.3. It consistently surpasses baseline LoRA (+2.3+2.3+ 2.3), LoRA-SAM (+1.0+1.0+ 1.0), and both BAR variants. These results highlight its superior generalization ability under distribution shift and limited supervision. We conjecture that the performance gap between SAM and EFlat-LoRA comes from EFlat-LoRA eliminating the mutual interference between perturbations in the two low-rank subspaces.
Fine-tuning with RoBERTa-large. We apply EFlat-LoRA to finetune RoBERTa-large. Our implementation is inspired from (Hu et al. 2022). The hyperparameters are chosen the same as provided in its GitHub repository. We report results on the development set, as done in previous work (Hu et al. 2022), since the test set is not publicly available. The results can be found in Table 2. we observe that EFlat-LoRA achieves the highest scores on all datasets, and achieves the highest accuracy on average over these datasets. Specifically, on average over these datasets, EFlat-LoRA surpasses standard LoRA with a margin of 1.01.01.0. Additionally, EFlat-LoRA even achieve higher than full fine-tuning on some datasets. This superior performance may be attributed to overfitting in full fine-tuning, where optimizing all model parameters can lead to overfitting on the training data, thus reducing the model’s generalization to the test set. This effect is particularly pronounced on small datasets, such as MRPC, which contains only 3.7k training data.
Fine-tuning with GPT-2. Having established that Flat-LoRA is an effective alternative to LoRA in NLU tasks, we now turn to explore whether EFlat-LoRA can further improve upon LoRA in the context of NLG models, such as GPT-2 Medium and Large (Radford et al. 2019). To enable a direct comparison, we adopt the experimental setup of (Li and Liang 2021) with minimal deviation. Table 3 demonstrates the effectiveness of EFlat-LoRA on the E2E NLG Challenge (Novikova, Dušek, and Rieser 2017) with GPT-2 Medium and Large models. Compared with existing PEFT methods such as Adapter and LoRA, EFlat-LoRA consistently achieves superior performance across all metrics. Notably, it achieves this improvement without increasing the number of trainable parameters, maintaining the same efficiency as standard LoRA.
@@
Few-shot with CLIP. Recent advances in few-shot adaptation of Vision-Language Models (VLMs) have significantly enhanced their generalization capabilities, requiring only a few labeled samples for downstream tasks. CLIP-LoRA (Zanella and Ben Ayed 2024) explores the application of LoRA in this few-shot VLM setting. In our work, we also apply Flat-LoRA and EFlat-LoRA to VLMs to evaluate their effectiveness. For a fair comparison, our experimental setup follows that of CLIP-LoRA (Zanella and Ben Ayed 2024). We consider five datasets for fine-grained classification of satellite imagery (EuroSAT (Helber et al. 2019), Ox-fordPets (Parkhi et al. 2012), Flower102 (Nilsback and Zisserman 2008), Caltech101 (Fei-Fei, Fergus, and Perona 2004), DTD (Cimpoi et al. 2014)). These datasets offer a thorough benchmarking framework for evaluating few-shot visual classification tasks. Table 4 demonstrates that Flat-LoRA and EFlat-LoRA outperformed Adapter and LoRA in most settings. In the low-data regimes (1-shot and 4-shot), EFlat-LoRA shows clear advantages. These results highlight the effectiveness of EFlat-LoRA in improving generalization in few-shot adaptation of vision-language models.
@@
Flat-LoRA 72.6 92.8 82.8 94.5 54.9
EFlat-LoRA 78.3 92.8 81.0 93.9 54.6
@@
Flat-LoRA 90.0 93.1 94.9 95.6 65.7
EFlat-LoRA 87.6 91.1 94.0 95.6 65.0
@@
Flat-LoRA 92.2 93.4 98.5 96.5 72.7
EFlat-LoRA 91.6 91.5 98.1 96.6 71.9
Table 4: Detailed results for five datasets with CLIP-Adapter, CLIP-LoRA and EFlat-LoRA.
Fine-tuning with Qwen-VL-Chat. Qwen-VL-Chat (Bai et al. 2023) is a multimodal conversational large language model capable of understanding both images and text. We apply EFlat-LoRA to fine-tune Qwen-VL-Chat, following the same experimental setup as in (Zhou et al. 2024). Table 5 presents the results on the ScienceQA (Lu et al. 2022) and VizWiz (Gurari et al. 2018) datasets. The results in Table 5 demonstrate that the perturbation size $\rho$ significantly influences the performance of EFlat-LoRA when fine-tuning Qwen-VL-Chat. By tuning $\rho$, EFlat-LoRA adapts to different tasks, enabling improved generalization—achieving higher accuracy than LoRA. Specifically, a larger $\rho$ (e.g., $\rho=0.2$) yields the best accuracy on ScienceQA, while a smaller $\rho$ (e.g., $\rho=0.05$) performs better on VizWiz. This suggests that different tasks benefit from different levels of perturbation. Therefore, selecting an appropriate $\rho$ based on the task characteristics is crucial for achieving optimal fine-tuning performance on multimodal large language models.
@@
Table 5: EFlat-LoRA Fine-Tuning Results on Qwen-VL-Chat with different $\rho$.

### v2

@@ Introduction
FMLoRA 95.1±0.5 54.4±1.3 86.4±0.8 82.7±1.0 82.7±1.2 96.7±0.2 83.1
EFMLoRA 91.9±1.7 54.7±1.6 85.7±0.7 82.1±0.6 82.8±0.2 96.8±0.4 82.3
Table 1: Experiments on few-shot RoBERTa (355M). Results marked with * are taken from (li2024implicit).
To theoretically analyze the error of EFMLoRA, some necessary assumptions are listed below, all of which are common and standard when analyzing SAM optimization (du-2022-ESAM-ICLR) (zhuang-2022-GSAM-ICLR).
Assumption 1.
@@
This convexity assumption is reasonable in the fine-tuning stage, as the model is typically close to a local minimum and the loss landscape is approximately convex in a local neighborhood (jang2024lora).
Theorem 2.
[EMA perturbation approximate perturbation of SAM due to the convex of the loss landscape] Assume that during fine-tuning, the solution is already close to a local minimum and the local loss function is convex. Let the model weights at $i$-th iteration be $\mathbf{w}_{t}$. Under Assumptions 1, 2, and 3, let ${\rho_{t}}=\frac{{{\rho_{0}}}}{{\sqrt{t}}}$, the error between the sharpness calculated using the EMA perturbation ($S^{\text{EMA}}$) and that calculated using the original SAM perturbation ($S^{\text{SAM}}$) is bounded as follows:
$|\underbrace{\left[{L({\mathbf{w}_{t}}+{\bm{\hat{\varepsilon}}_{t}})-L({\mathbf{w}_{t}})}\right]}_{S^{\text{EMA}}}-\underbrace{\left[{L({\mathbf{w}_{t}}+{\bm{\tilde{\varepsilon}}_{t}})-L({\mathbf{w}_{t}})}\right]}_{S^{\text{SAM}}}|$ (19)
$\leq\left({\left({1+{{(1-\beta)}^{t-1}}}\right)\tau{\rho_{0}}+G+{\sigma^{2}}}\right)$
$\quad\quad\cdot\left({\left({1+{{(1-\beta)}^{t-1}}}\right){\rho_{0}}+\frac{{{\rho_{0}}}}{{\sqrt{t}}}}\right).$
@@
[…]
@@
EFMLoRA 96.3±0.2 92.7±0.1 89.3±0.6 91.6±0.1 94.8±0.1 91.5±0.4 90.7±0.1 68.0±1.2 89.4
Table 2: Experiments on finetuning RoBERTa (355M). Results marked with $\dagger$ are taken from (hu2022lora), and those with * are taken from (li2024implicit).
@@
GPT-2 M (FMLoRA) 0.35M 69.2 8.72 46.6 71.5 2.51
GPT-2 M (EFMLoRA) 0.35M 69.7 8.77 46.6 71.7 2.53
@@
GPT-2 L (FMLoRA) 0.77M 70.0 8.83 46.8 71.8 2.53
GPT-2 L (EFMLoRA) 0.77M 70.2 8.84 46.8 71.8 2.52
Table 3: GPT-2 medium (M) and large (L) with different adaptation methods on the E2E NLG Challenge. Results marked with $\dagger$ are taken from (hu2022lora).
@@
LoRA reduces the number of trainable parameters by decomposing weight updates as $\Delta\mathbf{W}\approx\mathbf{B}\mathbf{A}$, where $\mathbf{B}\in\mathbb{R}^{n\times r}$ and $\mathbf{A}\in\mathbb{R}^{r\times m}$ with $r\ll\min(n,m)$. Both FMLoRA and EFMLoRA retain this parameter efficiency:
$\text{P}_{\text{LoRA}}$ $=\text{P}_{\text{FMLoRA}}=\text{P}_{\text{EFMLoRA}}$ (20)
@@
However, FMLoRA and EFMLoRA introduce additional memory overhead. Specifically, FMLoRA temporarily stores the original values of $\mathbf{B}$ and $\mathbf{A}$, as well as the gradients of $\mathbf{A}$. The memory usage of FMLoRA is calibrated as follows:
@@
where $\text{M}_{\text{LoRA}}$ indicates the memory required by LoRA. The memory of EFMLoRA needs to maintain the EMA perturbation on $\mathbf{B}$ as follows:
$\text{M}_{\text{EFMLoRA}}=\text{M}_{\text{LoRA}}+O(2\times(nr+rm)).$ (22)
@@
For time complexity, suppose that the time complexity of optimizing the model with LoRA is $O(T)$, which mainly includes the time for forward and backward. Theoretically, the time complexity of FMLoRA is approximately as follows:
$\text{T}_{\text{FMLoRA}}\approx O(2T)=2\times\text{T}_{\text{LoRA}}.$ (23)
In contrast, the time complexity of EFMLoRA can be approximated as follows:
$\text{T}_{\text{EFMLoRA}}\approx O(T)=\text{T}_{\text{LoRA}}.$ (24)
We implement QR decomposition by Householder transformations, with time complexity of $O(r^{2}n)$ for an $r\times n$ matrix, e.g., $r$ is rank, $n$ is the input dimension in LORA.
@@
The best and second-best results are highlighted in bold and underline, respectively. Additional experimental details are provided in the supplementary file.
@@
Few-shot with RoBERTa-large. We first consider few-shot learning with EFMLoRA. Following the setup of (li2024implicit), we adopt RoBERTa-large—a 355M-parameter language model—as the backbone. The results in Table 1 show that FMLoRA outperforms all other methods with the highest average score (83.1), particularly excelling on SST-2, SNLI, and MNLI. EFMLoRA follows closely with an average score of 82.3. It consistently surpasses baseline LoRA (+2.3+2.3), LoRA-SAM (+1.0+1.0), and both BAR variants. These results highlight its superior generalization ability under distribution shift and limited supervision. We conjecture that the performance gap between SAM and EFMLoRA comes from EFMLoRA eliminating the mutual interference between perturbations in the two low-rank subspaces.
Fine-tuning with RoBERTa-large. We apply EFMLoRA to finetune RoBERTa-large. Our implementation follows (hu2022lora), using the same hyperparameters as those in its GitHub repository. The results can be found in Table 2. we observe that EFMLoRA achieves the highest scores on all datasets, and achieves the highest accuracy on average over these datasets. Specifically, on average over these datasets, EFMLoRA surpasses standard LoRA with a margin of $1.0$. Additionally, EFMLoRA even achieve better performance than full fine-tuning on some datasets. This superior performance may be attributed to overfitting in full fine-tuning, where optimizing all model parameters can lead to overfitting on the training data, thus reducing the model’s generalization to the test set. This effect is particularly pronounced on small datasets, such as MRPC, which contains only 3.7k training data.
Fine-tuning with GPT-2. Having shown that FMLoRA is effective for NLU tasks, we now explore whether EFMLoRA can improve LoRA in NLG models like GPT-2 Medium and Large (radford2019language). To enable a direct comparison, we adopt the experimental setup of (li2021prefix) with minimal deviation. Table 3 demonstrates the effectiveness of EFMLoRA on the E2E NLG Challenge (novikova2017e2e) with GPT-2 Medium and Large models. Compared with existing PEFT methods such as Adapter and LoRA, EFMLoRA consistently achieves superior performance across all metrics. Notably, it achieves this improvement without increasing the number of trainable parameters, maintaining the same efficiency as standard LoRA.
@@
Few-shot with CLIP. Recent advances in few-shot adaptation of Vision-Language Models (VLMs) have significantly enhanced their generalization. CLIP-LoRA (zanella2024low) explores the application of LoRA in this few-shot VLM setting. In our work, we also apply FMLoRA and EFMLoRA to VLMs to evaluate their effectiveness. For a fair comparison, our experimental setup follows that of CLIP-LoRA. We consider five datasets for fine-grained classification of satellite imagery (EuroSAT (helber2019eurosat), Ox-fordPets (parkhi2012cats), Flower102 (nilsback2008automated), Caltech101 (fei2004learning), DTD (cimpoi2014describing)). These datasets offer a thorough benchmarking framework for evaluating few-shot visual classification tasks. Table 4 demonstrates that FMLoRA and EFMLoRA outperformed Adapter and LoRA in most settings. In the low-data regimes (1-shot and 4-shot), EFMLoRA shows clear advantages. These results highlight the effectiveness of EFMLoRA in improving generalization in few-shot adaptation of vision-language models.
@@
FMLoRA 72.6 92.8 82.8 94.5 54.9
EFMLoRA 78.3 92.8 81.0 93.9 54.6
@@
FMLoRA 90.0 93.1 94.9 95.6 65.7
EFMLoRA 87.6 91.1 94.0 95.6 65.0
@@
FMLoRA 92.2 93.4 98.5 96.5 72.7
EFMLoRA 91.6 91.5 98.1 96.6 71.9
Table 4: Detailed results for five datasets with CLIP-Adapter, CLIP-LoRA and EFMLoRA.
Fine-tuning with Qwen-VL-Chat. Qwen-VL-Chat (Bai2023QwenVLAV) is a multimodal conversational large language model capable of understanding both images and text. We apply EFMLoRA to fine-tune Qwen-VL-Chat, following the same experimental setup as in (zhou2024empirical). Table 5 presents the results on the ScienceQA (lu2022learn) and VizWiz (gurari2018vizwiz) datasets. The results in Table 5 demonstrate that the perturbation size $\rho$ significantly influences the performance of EFMLoRA when fine-tuning Qwen-VL-Chat. By tuning $\rho$, EFMLoRA adapts to different tasks, enabling improved generalization—achieving higher accuracy than LoRA. Specifically, a larger $\rho$ (e.g., $\rho=0.2$) yields the best accuracy on ScienceQA, while a smaller $\rho$ (e.g., $\rho=0.05$) performs better on VizWiz. This suggests that different tasks benefit from different levels of perturbation. Therefore, selecting an appropriate $\rho$ based on the task characteristics is crucial for achieving optimal fine-tuning performance on multimodal large language models.
@@
Table 5: EFMLoRA Fine-Tuning Results on Qwen-VL-Chat with different $\rho$.
@@
$downarrow Memory\downarrow Time\downarrow$
LoRA 23.6 4.30 23.2 8.45
FMLoRA 24.0 9.10 23.6 17.47
EFMLoRA 24.0 4.80 23.2 10.00
Table 6: Runtime (Hour) and memory (GB) of LoRA, FMLoRA and EFMLoRA on fine-tuning GPT-2 Medium/Large.
Runtime and Memory Consumption
The results in Table 6 confirm the theoretical time complexity analysis. As expected, FMLoRA has approximately double the runtime of LoRA (2.1× on both GPT-2 Medium and Large), consistent with its theoretical complexity of $O(2T)$ due to two forward and backward passes for sharpness optimization. In contrast, EFMLoRA operates with near-LoRA efficiency, requiring only 1.1× and 1.2× more time on GPT-2 Medium and Large, respectively. This supports the theoretical claim that EFMLoRA maintains a time complexity close to $O(T)$ while benefiting from sharpness-aware optimization. In addition, EFMLoRA maintains a memory usage almost identical to that of LoRA, with only negligible increases (less than 0.4 GB across both model scales). These results demonstrate that EFMLoRA achieves near-LoRA efficiency in both memory and runtime.
Conclusion
In this work, we propose FMLoRA, a novel PEFT method that integrates sharpness-aware optimization into the LoRA framework to promote convergence toward flatter minima. We theoretically demonstrate that perturbations in the full parameter space can be equivalently represented within the low-rank subspace. To improve computational efficiency, we introduce EFMLoRA, which leverages an exponential moving average to approximate perturbations, significantly reducing runtime overhead while maintaining effectiveness. Extensive experiments across various large language and vision-language models demonstrate that EFMLoRA achieves comparable or even superior generalization performance to full fine-tuning and LoRA. Our results emphasize the importance of reducing sharpness to improve generalization in PEFT methods, offering valuable insights and practical tools for future research on the link between sharpness and generalization in LLMs and beyond.
A. Proofs
A.1 Proof of Eq.(10) and Eq.(11)
Proof.
we propose to approximate the unknown gradient $\nabla L_{\mathbf{W}}(\mathbf{W})$ using standard LoRA gradients, which can be computed in two ways:
$(1)\nabla L_{\mathbf{B}}$ $(\mathbf{W_{0}}+s\mathbf{BA})=s\nabla L_{\mathbf{W}}(\mathbf{W})\mathbf{A}^{\top}$
$\Rightarrow\quad\nabla L_{\mathbf{W}}(\mathbf{W})=\frac{1}{s}\nabla L_{\mathbf{B}}(\mathbf{W_{0}}+s\mathbf{BA})(\mathbf{A}^{\top})^{+},$ (25)
$(2)\nabla L_{\mathbf{A}}$ $(\mathbf{W_{0}}+s\mathbf{BA})=s\mathbf{B}^{\top}\nabla L_{\mathbf{W}}(\mathbf{W})$
$\Rightarrow\quad\nabla L_{\mathbf{W}}(\mathbf{W})=\frac{1}{s}(\mathbf{B}^{\top})^{+}\nabla L_{\mathbf{A}}(\mathbf{W_{0}}+s\mathbf{BA}),$ (26)
A.2 Proof of Theorem 1
Proof.
The update process of the Flat-LoRA is as follows:
${\tilde{\mathbf{x}}_{t}}={{\mathbf{x}}_{t}}+\rho\frac{1}{s}\frac{{{\mathbf{G}_{t}}}}{{\left\|{{\mathbf{G}_{t}}}\right\|}}{\mathbf{y}_{t}}^{+}$ $,\quad{\tilde{\mathbf{y}}_{t}}={{\mathbf{y}}_{t}}$ (27)
${\mathbf{g}_{{\tilde{\mathbf{x}}_{t}}}}={{\tilde{\mathbf{G}}}_{t}}\tilde{\mathbf{y}}_{t}$ $,\quad{\mathbf{g}_{{\tilde{\mathbf{y}}_{t}}}}={{\tilde{\mathbf{G}}}_{t}}^{\top}\tilde{\mathbf{x}}_{t}$
${{\mathbf{x}}_{t+1}}={{\mathbf{x}}_{t}}-\eta{\mathbf{g}_{{\tilde{\mathbf{x}}_{t}}}}$ $,\quad{{\mathbf{y}}_{t+1}}={{\mathbf{y}}_{t}}-\eta{\mathbf{g}_{{\tilde{\mathbf{y}}_{t}}}}$